OpenStudy (anonymous):

A ball is thrown upward. What is its initial vertical speed? The acceleration of gravity is 9.8 m/s^2 and maximum height is 6.1 m. Neglect air resistance. Answer in units of m/s.

OpenStudy (anonymous):

From the information given, the ball will come to rest at the maximum height (it cannot go past this point) Kinematic Equation: Vf^2 = Vi^2 +2a(Xf - Xi) 0 = Vi^2 + 2(-9.8)(6.1 - 0) 0 = Vi^2 - 119.56 119.56 = Vi^2 Vi = 10.9 m/s